Suppose X follows the uniform distribution from the range (-a,a) where a is greater than zero. Find the value of a such that P(X<1)=(1/3)=P(X>1)
1
Expert's answer
2020-05-11T14:33:30-0400

A continuous rv "X" is said to have a uniform distribution on the interval "[A,B]" if the pdf of "X" is


"f(x;A,B) = \begin{cases}
 {1\over B-A} &A\leq x\leq B\\
 0 &otherwise
\end{cases}"

Suppose "X" follows the uniform distribution from the range (-a,a) where a is greater than zero.


"f(x;-a,a) = \begin{cases}
 {1\over 2a} & -a\leq x\leq a\\
 0 &otherwise
\end{cases},\ a>0"

"P(X<-1)=\displaystyle\int_{-\infin}^{-1}f(x)dx=\displaystyle\int_{-a}^{-1}{1\over 2a}dx=""={1\over 2a}[x]\begin{matrix}
 -1 \\
 -a
\end{matrix}=-{1\over 2a}+{1\over 2},a\geq1"

"P(X<1)=\displaystyle\int_{-\infin}^{1}f(x)dx=\displaystyle\int_{-a}^{1}{1\over 2a}dx=""={1\over 2a}[x]\begin{matrix}
 1 \\
 -a
\end{matrix}={1\over 2a}+{1\over 2},a\geq1"

Given "P(X<-1)=\dfrac{1}{3}"


"P(X<-1)=-{1\over 2a}+{1\over 2}={1\over 3}""{1\over 2a}={1\over 6}""a=3"

"P(X>1)=1-P(X<1)=1-({1\over 2(3)}+{1\over 2})={1\over 2}-{1\over 6}={1\over 3}"

"a=3."